OpenStudy (anonymous):

Find equations of the tangent line and normal line to the curve at the given point.

OpenStudy (anonymous):

\[y=\frac{ \sqrt{x} }{ x+9 }\]

OpenStudy (anonymous):

Given point (1,0.1)

OpenStudy (welshfella):

first you nee to find the derivative y' . This will give you the slope of the line in terms of x.

OpenStudy (anonymous):

\[\frac{(x+9)(\frac{ 1 }{ 2 } x ^{-1/2})-(x ^{1/2})(1)}{ (x+9)^2}\]

OpenStudy (welshfella):

yes thats correct now plug in x = 1 to find the numerical value of slope at the point (1,0.1)

OpenStudy (welshfella):

1/2 * (1)^(-1/2) = 1/2 * 1 /1^ 1/2

OpenStudy (welshfella):

= 1/2

OpenStudy (anonymous):

is the slope 4/100?

OpenStudy (welshfella):

right or 0.04

OpenStudy (anonymous):

y-0.1=.04(x-1)

OpenStudy (anonymous):

y=.04x-.04+0.1

OpenStudy (welshfella):

yes thats the tangent

OpenStudy (anonymous):

y=.04x+.06

OpenStudy (welshfella):

no - i think you have the last number wrong

OpenStudy (anonymous):

ohh

OpenStudy (welshfella):

sorry you are right!! my bad arithmetic!

OpenStudy (anonymous):

ok xD

OpenStudy (welshfella):

Now the normal as a slope of -1 / 0.04

OpenStudy (anonymous):

is the normal line perpendicular?

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

y-.1=-25(x-1)

OpenStudy (anonymous):

y=-25x+25+.1

OpenStudy (anonymous):

y=-25x+25.1

OpenStudy (welshfella):

right

OpenStudy (anonymous):

Thank you so much! You were so helpful :)

OpenStudy (welshfella):

yw

OpenStudy (welshfella):

you did pretty well too!!

OpenStudy (anonymous):

aw thanks :)